### v1

# A Paradigm for Creative Ownership

A Paradigm for Creative Ownership
Tejaswi Polimetla  and  Katy Ilonka Gero Harvard UniversityCambridgeMassachusettsUSA tpolimetla@seas.harvard.edu
(April, 2025; May 2025; May 2025)
creative ownership, human-AI collaboration, design frameworks, authorship, co-creation
copyright: noneccs: Human-centered computing HCI theory, concepts and modelsccs: Social and professional topics Intellectual propertyccs: Human-centered computing Collaborative and social computing theory, concepts and paradigms
*Presented at the Navigating Generative AI Disclosure, Ownership, and Accountability in Co-Creative Domains Workshop at CHIWORK 2025.
1. Introduction
With the rise of generative AI tools, the determination of ownership in collaborative human-AI settings is a prominent challenge. While legal experts have approached ownership from intellectual property and copyright law, literature in the social sciences and humanities reveals how humans come to form feelings of ownership in the first place. At the same time, while these fields have many conceptualizations of ownership, there does not exist a comprehensive system to diagnose where feelings of ownership come from, why these sentiments arise, and how they can be influenced. In their 2003 paper (Pierce et al., 2003), Pierce et al. define psychological ownership as ”that state where an individual feels as though the target of ownership or a piece of that target is ‘theirs’.” In this paper, we will focus on a narrower definition revolving around creative ownership in which the target of ownership is a creative product or artifact that the individual in question had a role in creating - no matter how small or large. By understanding how humans form feelings of ownership in creative contexts, we hope to form better intuitions in a variety of ownership contexts, as well as find ways to support system designers in preserving human ownership and authenticity. In this paper, we are interested in achieving the following goals: 1) identifying the key criteria required for one to develop a sense of creative ownership over one’s work, and 2) developing an explanatory framework for understanding why different circumstances of creation might lead to different levels of ownership sentiment.
2. Literature Review
The concept of ownership has been studied widely across the humanities and social sciences, with each field having a unique approach to the subject matter (Pipes, 2000). This literature review focuses on the fields of philosophy and psychology as these two fields most readily discuss issues related to how humans develop feelings of ownership on an individual level. Other social sciences such as economics, politics, and sociology are concerned with broader system level concerns. While other humanistic fields such as art and literature do discuss individual level concerns, the arguments developed are often grounded in formal philosophical reasoning.
2.1. Philosophy
While there are many schools of philosophical thought that could be used to frame a discussion of ownership, two juxtaposing ones that encompass the duality of ownership related values are materialism and idealism. Materialist theories stem from notions of property as control over material entities, going as far as to stipulate that physical, material states are the ultimate determinants of reality, taking precedence over thought, consciousness, and abstract entities (Melnyk, 2012; Smart, 1963). On the contrary, idealism posits that something mental is the ultimate foundation of reality, and idealist theories of property and personhood are concerned with symbolic and mental conceptions of ownership (Guyer and Horstmann, 2015). This dualistic framing captures both the tangible and intangible elements of ownership.
One of the most fundamental materialist theories is Locke’s labor theory, which posits that “every man has a property in his own person,” and thereby goes on to argue that when one mixes their labor with natural resources, the resulting good becomes their property - evoking the embodiment theory of personhood (Locke, 1689; Radin, 1982). “Bundle of Rights” views hold ownership as a set of contractual obligations between people in relation to property (Waldron, 1985; Unknown, [n. d.]; Honoré, 1961). More broadly, these theories lend themselves as a basis for the construction of property systems as they tend to be grounded in tangible, enforceable, socially structured claims. In contrast, idealist theories tend to be much more abstract. For instance, Hegel’s ideas of ownership stem from the notion that the ”will” can be embodied in external entities, and that this embodiment is necessary for one’s actualization as a person cannot come to exist without both relation to and differentiation from the external environment (Radin, 1982). While the specifics of theories vary, the investment of one’s self, values, and identity as a means of developing feelings of ownership is a common theme that arises (Sartre, 1943; Émile Durkheim, 1957).
Although these views are contrasting, both reveal key features of ownership such as the investment of labor and the reflection of identity that contribute to a broader conception and intuition of how feelings of ownership are developed. It is evident that both the material and immaterial have appealed to human sensibilities throughout history.
2.2. Psychology
In the field of psychology, there have been numerous theoretical propositions and empirical studies attempting to explain the formation of psychological ownership. Several scholars have created frameworks based on decades of psychological research that capture key themes that have emerged time and again such as effectance and control of possessions (Furby, 1978; White, 1959; McClelland, 1951), positive affect (Furby, 1978), and symbolic meaning and personhood (Rochberg-Halton, 1980). These frameworks span a range of formulations ranging from Targets-Antecedents-Consequences-Interventions (Lyu and [Additional Authors, 2023) to corrective dual-process models (Morewedge, 2021), among others (Boyer, 2023; Pierce et al., 2001). Some of the major themes found across frameworks include responsibility, accountability, identity, self-efficacy, belongingness, control, self-congruity, psychological closeness, object-knowledge, self-investment, and rights over the object. While these frameworks capture numerous dimensions of psychological ownership, they fall short in creative contexts, where ownership by creation begins to deviate from other forms of ownership. Research on the self-creation effect illustrates how creating something oneself can lead to stronger object valuation and a more profound sense of ownership - aspects that are often overlooked by traditional frameworks of ownership (Koster et al., 2015; Brunneder and Dholakia, 2018). Therefore, we draw upon existing frameworks and approaches to produce a framework that is more streamlined for creative contexts.
In addition to these framework based approaches, there has also been much empirical work to understand the emergence of psychological ownership under particular circumstances. In a study of ePortfolio creation in online learning environments, researchers found that psychological ownership is correlated with perceived control of content, planning, personal data and access rights (Buchem, 2012). Levene et al. examined the effects of creation and intent on ownership judged and found that the effects of creation hold even when controlling for other factors. They also showed that successful and intentional creations are ascribed more ownership than unsuccessful or unintentional creations, and that creation is ascribed more ownership than the equivalent labor (Levene and Friedman, 2015). A study of locus of control and psychological ownership found internal locus of control to be positively related to both effectance and self-identity motives (McIntyre et al., 2009).
2.3. Other Areas
In addition to philosophy and psychology, there are a plethora of fields that discuss the emergence of ownership sentiments, specifically in creative contexts. Linguistics literature outlines how people speak about possession and ownership, highlighting how kinship and whole-part relations are indicative of closer links compared to possessor-possessee relations (Aikhenvald, 2012). In her experience as a community dance practitioner, Rosemary Lee details how she focuses on releasing the potential of dancers, with an attention to subtle creative choices in her process (Lee, 2008). In a study of psychological ownership in creative work, Rouse identifies differences in task type (managerial, conceptual, production), task target (abstract, concrete), and task coordination (independent, autonomous) (Rouse, 2013). When measuring ownership sentiments in academic writing contexts, Nicholes found that ownership sentiments tend to be significantly higher in creative genres as opposed to academic genres, and higher for argumentative papers as compared to research papers (Nicholes, 2015). Struggle is another theme that appears, with productive and creative struggle arising as both educational and design goals (Fox, 2023; Zhou and Sterman, 2023).
3. Framework
Person Process System
Embodiment - How much do you feel that the finished product embodies your values, personality, and identity? Control - How much control did you have over the creative decisions that shaped the final product? Production - How closely did you physically work with final product?
Occupancy - How often do you feel preoccupied with thinking about the product? Intentionality - How intentional were you about the creative decisions that you made? Abstraction - How closely did you conceptually work with the final product?
Recognition - Do you receive a lot of external recognition for your work? Effort - How much physical and/or mental effort did you put into this work? Interdependence - How independent were you during the creative process?
The proposed framework incorporates the material and immaterial components of ownership, tying together elements of individual values and personhood, the creative process, and the systems in which these individuals and processes are embedded. Personhood in a creative context is conceptualized as how the self relates to the creative outcome. How does the creative product embody one’s identity, self and values, how consistently does it occupy the mind of the creator, and to what extent do others recognize the creator (and their personhood) for their work? The creative process is the series of steps by which the creator’s personhood is realized in the creation, and is characterized by the intentionality of the creator, the amount of control they have over the process, and the amount of effort - both mental and physical - that goes into the realization. The system refers to the environment in which this creative process is enacted and its dimensions are borrowed from Rouse’s findings (Rouse, 2013). Abstraction refers to the creator’s role in conceptualizing the work, while production relates to their responsibility for its materialization. Interdependence reflects the creator’s reliance on others or tools to complete the creation. While process and system dimensions may be correlated at times, this is not always the case as shown in Appendix A.
We choose to visualize this framework using three concentric circles, each representing one of the elements listed above, and each element being composed of sub-dimensions. An increase in one dimension leads to an increase not only in the size of its corresponding circle, but in the size of the totality of the three circles altogether. Additionally, this visualization refrains from centering on any one element. Instead, precedence is largely context dependent, and the center-most circle can be thought of as the limiting factor in developing feelings of ownership. There is a threshold slider which is meant to indicate how different individuals may have different thresholds for developing feelings of ownership. The visualization can be explored using this interactive webtool: Creative Ownership Test.
A simple example of the framework’s application is the comparison between a master painter, like Picasso or Van Gogh, and a hobbyist art student (visualized in Appendix C). The master painter is widely recognized, with their work embodying their personhood. They invested significant time, intentionality, control, and effort into their art, working directly with the final product and maintaining independence throughout the creative process. In contrast, the hobbyist art student, following a paint-along workshop, is unlikely to view the finished painting as embodying their values. They have little control or intentionality in the outcome, exert moderate effort, and are highly reliant on the instructor. While they closely work with the final product, they play a minimal role in its conceptualization. From the figure, we can see that their main ownership “axis” is that of being part of a system of creation—they enact a small but critical part of the painting creation. Refer to Appendix B for more examples.
In the context of human-AI collaboration, we can think about how the use of AI systems might diminish feelings of ownership by impacting one or more dimensions of this framework. For instance, an artist who uses generative AI to create a piece of digital art as opposed to a more traditional medium such as painting is not as directly involved in the material production of the painting even if they feel high levels of the other dimensions such as embodiment, occupancy, or even effort. Onlookers might attribute the finished work to the AI system, thereby reducing external recognition. In collaborative contexts, this framework may be applied to each of the creators involved in creating the finished work to understand attributions of ownership.
4. Future Work
The next steps for this work involve conducting a user study with creative practitioners and professionals to understand how closely the proposed framework aligns with the experiences of creatives in a range of fields. To do so, we will ask practitioners how strongly they feel ownership sentiments over a creative artifact that they have produced before walking them through the questions on the framework. We hope to understand whether reported feelings of ownership align with the results of the framework, whether some elements are more important than others or if there are any missing elements, and whether there are differences between various fields of practice. Following validation of the framework, we envision this as being a useful research tool for studying ownership in human-AI collaboration contexts.
Acknowledgements.
Heartfelt gratitude to Ami Mehta, Samuel Budnyk, Victor Arul, and Flora Saki Giordani for generously sharing their expertise and insights as artists and creators. Their contributions have been invaluable to this work.
Appendix A Demonstrating the Difference Between Process and System
This example will show that the process dimensions (control, intentionality, and effort) and the system dimensions (production, abstraction, and interdependence) are not always necessarily in alignment, even if it may appear to be so more generally. Imagine an individual who just purchased a paint-by-numbers kit and is completing the project at home. This individual does not have much control or intentionality over the creative decisions impacting the final painting, and may have put in little to moderate effort. In line with this, their score for conceptualization would be low as well as they did not directly conceive of the final outcome. However, their score for production would be very high as they directly worked with the final material outcome of the painting. Therefore, while process dimensions and system dimensions may be positively correlated in some scenarios, this is not always necessarily the case.
Appendix B Example Applications of the Framework
B.1. The Interior Designer
The interior designer, representing commissioned creatives more broadly, typically reflects the client’s preferences and values in their design choices even if the client may expect some aesthetic input from the designer. The designer invests significant time and receives moderate recognition from the client, firm, and potential guests. They have substantial control over the process and exert effort, but the final intentionality of the space depends on the client. Although they work closely with the final product and conceptualize the design, they collaborate with the client and other stakeholders to execute the vision, with their ownership primarily stemming from enacting the creative process.
B.2. The Community Arts Practitioner
The community arts practitioner might be an individual who has a grand creative vision and loops in their community to make it happen. Their ideas likely embody their values along with those of their community, they likely spend a lot of time thinking about the idea, and their community recognizes them for this piece of (generally public) artwork. They have a high degree of control over the creative process - the choices that go into deciding how to execute the piece - but when it comes to the actual execution itself, they are highly dependent on the community members to carry out the vision.
B.3. A Patron of the Arts
This is someone who has a grand vision for a piece of artwork and commissions an artist to execute it. The final artwork is intended to embody this person’s ideals, and they spend a considerable amount thinking about it, and many in art circles recognize them for their patronage. While they have a high degree of intentionality, they must cede some control to the artist, and they expend a moderate amount of effort. The artist is entirely responsible for the production, and the patron is considerably dependent on the artist for executing the vision, and must cede some of the conceptualization to the artist.
Appendix C Creative Ownership Framework Applications
Figure 1. Ownership Framework for a Master Artist
Figure 2. Ownership Framework for an Amateur Artist